OpenStudy (anonymous):

Simplify the square root of 12.

ganeshie8 (ganeshie8):

\(\large \sqrt{12} = \sqrt{4 \times 3} = \sqrt{4} \times \sqrt{3}\)

ganeshie8 (ganeshie8):

whats the sqrt(4) ?

OpenStudy (anonymous):

Okay thank you. I just like to over the lessons I learned.

OpenStudy (anonymous):

2*2

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

\(\large \sqrt{12} = \sqrt{4 \times 3} = \sqrt{4} \times \sqrt{3} = \sqrt{2*2} \times \sqrt{3} = 2\sqrt{3}\)

ganeshie8 (ganeshie8):

see if that looks okay :)

OpenStudy (anonymous):

4*6?

ganeshie8 (ganeshie8):

nope. the simplified form is just this : \(\large 2 \sqrt{3}\)

OpenStudy (anonymous):

Oh I see now.

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

Thank you so much @ganeshie8

ganeshie8 (ganeshie8):

np... u wlc :)